# Convergence of spectral discretization for the flow of diffeomorphisms

### v1

@@ Abstract
The Large Deformation Diffeomorphic Metric Mapping (LDDMM) or flow of diffeomorphism is a classical framework in the field of shape spaces and is widely applied in mathematical imaging and computational anatomy. Essentially, it equips a group of diffeomorphisms with a right-invariant Riemannian metric, which allows to compute (Riemannian) distances or interpolations between different deformations. The associated Euler–Lagrange equation of shortest interpolation paths is one of the standard examples of a partial differential equation that can be approached with Lie group theory (by interpreting it as a geodesic ordinary differential equation on the Lie group of diffeomorphisms). The particular group $\mathcal{D}^{m}$ of Sobolev diffeomorphisms is by now sufficiently understood to allow the analysis of geodesics and their numerical approximation. We prove convergence of a widely used Fourier-type space discretization of the geodesic equation. It is based on a new regularity estimate: We prove that geodesics in $\mathcal{D}^{m}$ preserve any higher order Sobolev regularity of their initial velocity.
@@ 1 Introduction and main results
There are a number of partial differential equations (PDEs) that can be interpreted as a geodesic equation (the Euler–Lagrange equation satisfied by locally shortest paths) on an infinite-dimensional Lie group. This viewpoint started with the seminal work by Arnol’d in the 1960s on hydrodynamics, who for instance interpreted the Euler equations of inviscid incompressible fluid flow as the geodesic equations on the Lie group of volume-preserving diffeomorphisms endowed with a right-invariant $L^{2}$-metric. The great advantage of such a viewpoint is that it makes these PDEs amenable to an analysis via ordinary differential equation (ODE) techniques. It is also exploited for numerics, e.g. by devising efficient, structure-preserving solvers based on Hamiltonian system integrators. Other examples of PDEs that fit into this framework include Burgers’ equation, the Camassa-Holm equation, and the KdV equation, see [MeP10, KW09] and the references therein.
@@
for $\mathcal{L}$ a self-adjoint differential operator such as $\mathcal{L}=(1-\Delta)^{m}$ with $m\geq 0$ ($t$ denotes time, and the dot time differentiation). The close similarity to the Euler equations becomes apparent for $m=0$, in which case the equation reads $\dot{v}_{t}+\mathrm{div}(v_{t}\otimes v_{t})+\nabla\frac{|v_{t}|^{2}}{2}=0$ (the incompressible Euler equations just differ by the additional incompressibility constraint $\mathrm{div}v_{t}=0$ and the replacement of the internal energy $\frac{|v_{t}|^{2}}{2}$ by the pressure $p$). This EPDiff equation occurs as the Euler–Lagrange equation when trying to deform (or rather transport) a given image into another one by a time-dependent velocity field $v_{t}$ with least possible energy $\int_{0}^{1}\langle\mathcal{L}v_{t},v_{t}\rangle\,\mathrm{d}t$. Therefore is is used a lot in computational anatomy, where medical images of a patient have to be mapped to an annotated template image and where this framework for dealing with deformations is known as Large Deformation Diffeomorphic Metric Mapping (LDDMM).
This EPDiff equation actually turns out to be the geodesic equation on the group $\mathcal{D}^{m}$ of diffeomorphisms (e.g. of the unit cube) of Sobolev regularity $m$, endowed with a right-invariant Riemannian Sobolev metric. A geodesic (i.e. a locally shortest path) $t\mapsto\phi_{t}$ of diffeomorphisms can be written as the so-called flow of a velocity field $v_{t}$, i.e. as solution of
@@
(in the language of fluid mechanics, $\phi_{t}$ describes the motion in Lagrangian coordinates, while $v_{t}$ is the Eulerian description of the motion), and this velocity field $v_{t}$ satisfies the EPDiff equation (where $\mathcal{L}$ is related to the employed Riemannian metric).
The EPDiff equation represents a typical model setting for applying Lie group techniques to PDEs, see e.g. [TY15, MeP10]. One of the reasons is that by now the group of Sobolev diffeomorphisms and its geometric properties are quite well understood [MeP10, BV17, GRRW23] and that the numerical implementation is straightforward. In particular, (1) lends itself to a space discretization by truncated Fourier series: For instance, the authors of [ZF19] propose a scheme in which $\rho_{t}$ and $v_{t}$ are approximated by bandlimited functions $P_{t}$ and $V_{t}$ (numerically represented by their finite Fourier series) and both right-hand sides in (1) are essentially just truncated in Fourier space to satisfy the band limit, yielding a highly efficient code. Their intuition is that the smoothing operator $\mathcal{R}$ produces an effective bandlimit anyway.
@@
Under standard conditions on $\mathcal{D}^{m}$ and its Riemannian metric, if the initial velocity $v_{0}$ has Sobolev regularity $m+k$ for $k\geq 1$, the numerical approximation of bandlimiting the right-hand sides in (1) converges to the true solution as the bandlimit $R$ tends to infinity. Moreover, for $k\geq 2$ the error tends to zero at rate $R^{1-k}$.
(Note that numerically the convergence order seems to be better by one, so generically the convergence might be better.) This shows that the intuition of [ZF19] may be only partly correct; the smoothing operator $\mathcal{R}$ alone might not produce a strong enough bandlimiting for the approximation to converge. (In fact, [ZF19] additionally replaces differentiation by finite differences, but this only slightly simplifies the implementation.) Following the theme that the underlying Lie group structure allows to use ODE techniques, the convergence is proven via Gronwall’s inequality. This is made possible by a new regularity result for geodesics in $\mathcal{D}^{m}$, theorem 11, a short form of which is the following.
@@
Under standard conditions on $\mathcal{D}^{m}$ and its Riemannian metric, if the initial velocity $v_{0}$ has Sobolev regularity $m+k$ for $k\geq 1$, then so does the velocity $v_{t}$ for all $t$.
This result builds upon the knowledge acquired during recent years on $\mathcal{D}^{m}$, in particular the well-posedness of geodesics, the existence of shortest geodesics, and the rigorous derivation of the geodesic equation [MeP10, BV17, GRRW23]. It confirms a conjecture of [MeP10] who could only show preservation of Sobolev regularity $m+k$ for $k>m+\frac{d}{2}$ in space dimension $d$.
@@
Finally note that LDDMM and the EPDiff equation are also frequently employed with alternative smoothing operators $\mathcal{R}$ such as convolution with a Gaussian. None of the shown analysis applies to these settings, since almost nothing is so far known about the geometry of the associated group of diffeomorphisms: Is it a manifold, and what (Banach) space $X$ is it modelled over? Is the right-invariant Riemannian metric induced by $\mathcal{L}=\mathcal{R}^{-1}$ smooth so that geodesics are well-defined? Our numerical analysis further makes use of the following properties: Is $X$ a Banach algebra so that one can make sense of the quadratic terms in (1)? Does an estimate of the form $\|(\mathcal{L}v)Dv\|_{(X^{+})^{\prime}}\leq\|v\|_{X}^{2}$ hold with $X^{+}=\{v\in X\,|\,Dv\in X^{d}\}$ and $(X^{+})^{\prime}$ its dual space? Do geodesics preserve the property $v_{t}\in X^{+}$ or even higher differentiability?
@@
Below we briefly introduce some notation employed throughout. We will work on the $d$-dimensional flat torus $\mathbb{T}^{d}$ (in applications $d\in\{2,3\}$), which will be identified with $[0,1)^{d}$ with periodic boundary conditions. The single stroke norms ||$\cdot$|| ⋅ | and $|\cdot|_{\infty}$ denote the Euclidean $\ell^{2}$- and the $\ell^{\infty}$-norm on finite-dimensional vectors or tensors. The identity matrix is $\mathrm{I}$, the identity operator on a function space is denoted by ${\mathcal{I}}$, and the identity function by $\mathrm{id}$. If we write $\mathrm{id}:\mathbb{T}^{d}\to\mathbb{R}^{d}$, then we identify $\mathbb{T}^{d}$ with $[0,1)^{d}$. The adjoint of a linear operator $P$ is written as $P^{*}$.
We will employ the following function spaces: $H^{l}(\mathbb{T}^{d})$ denotes the Sobolev space of scalar functions on the torus with square-integrable weak derivatives up to order $l$. For simplicity we will only work with integer orders; the special case $l=0$ refers to square-integrable Lebesgue functions. For $\mathbb{R}^{d}$-valued functions we use the notation $H^{l}(\mathbb{T}^{d};\mathbb{R}^{d})$. The notation $\|\cdot\|_{H^{l}}$ then refers to any one of the equivalent $H^{l}$-norms (for vector-valued functions simply taken componentwise). In contrast, $\vvvert\cdot\vvvert_{H^{m}}$ corresponds to one particular $H^{m}$-norm (the one chosen when defining a metric on the space of Sobolev diffeomorphisms). The corresponding dual spaces are denoted $H^{-l}(\mathbb{T}^{d})$ and $H^{-l}(\mathbb{T}^{d};\mathbb{R}^{d})$, respectively. Note that the diffeomorphisms will actually be functions in $H^{m}(\mathbb{T}^{d};\mathbb{T}^{d})$ (so domain and codomain are the torus); this simply means that the diffeomorphism is in $H^{m}$ if restricted to any simply connected neighbourhood and expressed in local coordinates of domain and co-domain (cf. [Tay23, § 4.3]; the local coordinates are obtained by the identification of $\mathbb{T}^{d}$ with $\mathbb{R}^{d}$ modulo the integer lattice). The Sobolev space of functions with $l$ essentially bounded derivatives is denoted by $W^{l,\infty}$ (domain and codomain are indicated as for $H^{l}$) with norm $\|\cdot\|_{W^{l,\infty}}$, where the special case $l=0$ indicates essentially bounded Lebesgue functions. $C^{n}$ denotes $n$ times continuously differentiable functions (again domain and codomain will be indicated as for $H^{l}$), with $C^{\infty}$ representing infinitely often differentiable functions. Finally, we will employ the Bochner space $L^{1}([0,1];H^{m}(\mathbb{T}^{d};\mathbb{R}^{d}))$ with norm $\|\cdot\|_{L^{1}([0,1];H^{m})}$ of absolutely integrable paths in $H^{m}(\mathbb{T}^{d};\mathbb{R}^{d})$.
@@ 2 The Riemannian manifold of Sobolev diffeomorphisms
It is well-known that $\mathcal{D}_{\mathrm{id}}^{m}$ is a differential ($C^{\infty}$) manifold modelled over the Hilbert space $H^{m}(\mathbb{T}^{d};\mathbb{R}^{d})$ [MeP10]. In fact, this can be seen by interpreting $\mathcal{D}_{\mathrm{id}}^{m}$ as the qotient space $\widetilde{\mathcal{D}}_{\mathrm{id}}^{m}/\mathbb{Z}^{d}$ for
@@
Since $\mathcal{D}_{\mathrm{id}}^{m}$ (and thus also $\mathcal{D}^{m}$) is a group as well as a manifold, it behaves in some aspects like a Lie group and is often formally treated like one. However, inversion and left-multiplication are not smooth (not even differentiable), which is typically required for a Lie group. In fact, denoting the group of homeomorphisms and $C^{\infty}$-diffeomorphisms on the torus by $\mathrm{Hom}$ and $\mathrm{Diff}^{\infty}$, respectively, by [Bou72, III, Exerc. §4, ¶7] there cannot exist a Banach Lie group $G$ such that $\mathrm{Diff}^{\infty}\subset G\subset\mathrm{Hom}$ and the inclusions $\mathrm{Diff}^{\infty}\hookrightarrow G\hookrightarrow\mathrm{Hom}$ are continuous homomorphisms. In more detail, one can construct diffeomorphisms $\phi\in\mathrm{Diff}^{\infty}$ that are arbitrarily close to $\mathrm{id}$, but do not belong to any one-parameter subgroup of $\mathrm{Diff}^{\infty}$; thus they do not belong to a one-parameter subgroup of $G$ either. This, however, contradicts the local surjectivity of the Lie exponential map in Banach Lie groups.
@@
Let $(\!(\cdot,\cdot)\!)_{H^{m}}$ be an inner product on $H^{m}(\mathbb{T}^{d};\mathbb{R}^{d})$ with induced norm $\vvvert\cdot\vvvert_{H^{m}}$, then
@@
is the induced right-invariant Riemannian Sobolev metric on $\mathcal{D}_{\mathrm{id}}^{m}$.
@@
The geodesic equation is the optimality condition for the minimization of $E$ and thus a second order ODE on the manifold $\mathcal{D}_{\mathrm{id}}^{m}$ (which is a PDE on $[0,1]\times\mathbb{T}^{d}$). On honest Lie groups with right-invariant metric, the geodesic ODE is invariant under right action of the group on itself so that by Noether’s Theorem it can be reduced to a first order ODE. One can apply the same argument to formally derive such a first order geodesic equation on $\mathcal{D}_{\mathrm{id}}^{m}$: Let $\phi$ be a geodesic path with velocity $v$ and let $\eta\in C^{\infty}([0,1]\times\mathbb{T}^{d})$ be an infinitesimal perturbation with $\eta_{0}=\eta_{1}=0$. Defining $\phi^{\varepsilon}_{t}=(\mathrm{id}+\varepsilon\eta_{t})\circ\phi_{t}$, its velocity is given by $v^{\varepsilon}_{t}=\dot{\phi}^{\varepsilon}_{t}\circ(\phi^{\varepsilon}_{t})^{-1}=\left[(\mathrm{I}+\varepsilon D\eta_{t})v_{t}+\varepsilon\dot{\eta}_{t}\right]\circ(\mathrm{id}+\varepsilon\eta_{t})^{-1}$. Assuming sufficient differentiability, the optimality conditions read
@@
This first order ODE is known as the EPDiff or Euler–Poincaré equation [TY15]; its integral yields $\rho$ or equivalently $v$, from wich then the geodesic path $\phi$ can be calculated as the flow. Note that the derivation was merely formal since along the way we illegally differentiated the group product of $\mathcal{D}_{\mathrm{id}}^{m}$ (e.g. $\mathrm{Ad}_{\phi}$ with $\phi\in\mathcal{D}_{\mathrm{id}}^{m}$ maps $H^{m}(\mathbb{T}^{d};\mathbb{R}^{d})$ only into $H^{m-1}(\mathbb{T}^{d};\mathbb{R}^{d})$). As a consequence, due to $v_{t}\in H^{m}(\mathbb{T}^{d};\mathbb{R}^{d})$ and $\rho_{t}\in H^{-m}(\mathbb{T}^{d};\mathbb{R}^{d})$, the right-hand side of (6) only lies in $H^{-m-1}(\mathbb{T}^{d};\mathbb{R}^{d})$, and the equation may not be well-defined.
@@
    paths of finite energy in $\mathcal{D}_{\mathrm{id}}^{m}$ exist between any two elements of $\mathcal{D}_{\mathrm{id}}^{m}$ (essentially, for $\phi_{0}$ and $\phi_{1}$ close enough, the linear interpolation $\phi_{t}=t\phi_{1}+(1-t)\phi_{0}$ has finite energy),
@@
    for the inner product $(\!(v,w)\!)_{H^{m}}=\langle\mathcal{L}v,w\rangle$ with $\mathcal{L}=(1+\Delta)^{m}$ or a more general differential operator of order $2m$ with smooth coefficients the metric is smooth [MeP10, Thm. 4.1] (this is even claimed for any elliptic invertible operator $\mathcal{L}\in\mathrm{OPS}_{1,0}^{2m}$ of order $2m$), thus the geodesic ODE is locally uniquely solvable by classical differential geometry and ODE theory [Kli95, § 1.6], its solution depends smoothly on the initial conditions, and by the Inverse Function Theorem the Riemannian exponential map is a local diffeomorphism,

### v2

@@ Abstract
The Large Deformation Diffeomorphic Metric Mapping (LDDMM) or flow of diffeomorphism is a classical framework in the field of shape spaces and is widely applied in mathematical imaging and computational anatomy. Essentially, it equips a group of diffeomorphisms with a right-invariant Riemannian metric, which allows to compute (Riemannian) distances or interpolations between different deformations. The associated Euler–Lagrange equation of shortest interpolation paths is one of the standard examples of a partial differential equation that can be approached with Lie group theory (by interpreting it as a geodesic ordinary differential equation on the Lie group of diffeomorphisms). The particular group $\mathcal{D}^{m}$ of Sobolev diffeomorphisms is by now sufficiently understood to allow the analysis of geodesics and their numerical approximation. We prove convergence of a widely used Fourier-type space discretization of the geodesic equation. It is based on a regularity estimate, for which we also provide a new proof: Geodesics in $\mathcal{D}^{m}$ preserve any higher order Sobolev regularity of their initial velocity.
@@ 1 Introduction and main results
There are a number of partial differential equations (PDEs) that can be interpreted as a geodesic equation (the Euler–Lagrange equation satisfied by locally shortest paths) on an infinite-dimensional Lie group. This viewpoint started with the seminal work by Arnol’d in the 1960s on hydrodynamics [ARN66], who for instance interpreted the Euler equations of inviscid incompressible fluid flow as the geodesic equations on the Lie group of volume-preserving diffeomorphisms endowed with a right-invariant $L^{2}$-metric. The great advantage of such a viewpoint is that it often makes these PDEs amenable to an analysis via ordinary differential equation (ODE) techniques (corresponding sufficient conditions are e.g. given in [KOL17]). It is also exploited for numerics, e.g. by devising efficient, structure-preserving solvers based on Hamiltonian system integrators. Other examples of PDEs that fit into this framework include Burgers’ equation, the Camassa-Holm equation, and the KdV equation, see [MP10, KW09, KOL17] and the references therein.
@@
for $\mathcal{L}$ a self-adjoint differential operator such as $\mathcal{L}=(1-\Delta)^{m}$ with $m\geq 0$ ($t$ denotes time, the dot time differentiation, and $\mathcal{R}=\mathcal{L}^{-1}$). The close similarity to the Euler equations becomes apparent for $m=0$, in which case the equation reads $\dot{v}_{t}+\mathrm{div}(v_{t}\otimes v_{t})+\nabla\frac{|v_{t}|^{2}}{2}=0$ (the incompressible Euler equations just differ by the additional incompressibility constraint $\mathrm{div}v_{t}=0$ and the replacement of the internal energy $\frac{|v_{t}|^{2}}{2}$ by the pressure $p$). This EPDiff equation occurs as the Euler–Lagrange equation when trying to deform (or rather transport) a given image into another one by a time-dependent velocity field $v_{t}$ with least possible energy $\int_{0}^{1}\langle\mathcal{L}v_{t},v_{t}\rangle\,\mathrm{d}t$. Therefore it is used a lot in computational anatomy, where medical images of a patient have to be mapped to an annotated template image and where this framework for dealing with deformations is known as Large Deformation Diffeomorphic Metric Mapping (LDDMM).
This EPDiff equation actually turns out to be the geodesic equation on the group $\mathcal{D}^{m}$ of diffeomorphisms (e.g. of the unit cube) of Sobolev regularity $m$, endowed with a right-invariant Riemannian Sobolev metric: A geodesic (i.e. a locally shortest path) $t\mapsto\phi_{t}$ of diffeomorphisms can be written as the so-called flow of a velocity field $v_{t}$, i.e. as solution of
@@
(in the language of fluid mechanics, $\phi_{t}$ describes the motion in Lagrangian coordinates, while $v_{t}$ is the Eulerian description of the motion), and this velocity field $v_{t}$ satisfies the EPDiff equation (where $\mathcal{L}$ is related to the employed Riemannian metric $g$ via $g_{\mathrm{id}}(v,w)=\langle\mathcal{L}v,w\rangle$).
The EPDiff equation represents a typical model setting for applying Lie group techniques to PDEs, see e.g. [TY15, MP10]. One of the reasons is that by now the group of Sobolev diffeomorphisms and its geometric properties are quite well understood [MP10, BV17, GRR+23] and that the numerical implementation is straightforward. In particular, (1) lends itself to a space discretization by truncated Fourier series: For instance, the authors of [ZF19] propose a scheme on the $d$-dimensional torus $\mathbb{T}^{d}$ in which $\rho_{t}$ and $v_{t}$ are approximated by bandlimited functions $P_{t}$ and $V_{t}$ (numerically represented by their finite Fourier series) and both right-hand sides in (1) are essentially just truncated in Fourier space to satisfy the band limit, yielding a highly efficient code. Their intuition is that the smoothing operator $\mathcal{R}$ produces an effective bandlimit anyway.
@@
Let $m>1+\frac{d}{2}$ and the metric on $\mathcal{D}^{m}$ be induced by a Fourier multiplier $\mathcal{L}$. If the initial velocity $v_{0}$ has Sobolev regularity $m+k$ for $k\geq 1$, the numerical approximation of bandlimiting the right-hand sides in (1) converges to the true solution as the bandlimit $R$ tends to infinity. Moreover, for $k\geq 2$ the error tends to zero at rate $R^{1-k}$.
(Note that numerically the convergence order seems to be better by one, so generically the convergence might be better.) This shows that the earlier-mentioned intuition may be only partly correct: The smoothing operator $\mathcal{R}$ alone might not produce a strong enough bandlimiting for the approximation to converge, but an additional, stronger smoothing of the initial condition might potentially be required. (In fact, [ZF19] additionally replaces differentiation by finite differences, but this only slightly simplifies the implementation.) Following the theme that the underlying Lie group structure allows to use ODE techniques, the convergence is proven via Gronwall’s inequality. This is made possible by a regularity result for geodesics in $\mathcal{D}^{m}$, sometimes known as “no-loss-no-gain (of regularity)” [BHM25, Cor. 7.6], [BRU17, Sec. 5.2], the analysis of which started with [EM70]. In theorem 15 we prove a version slightly stronger than existing ones, a short form of which is the following.
@@
Let $m>1+\frac{d}{2}$ and the metric on $\mathcal{D}^{m}$ be strong, twice differentiable, and induced by some $\mathcal{L}=B^{*}B+\bar{\mathcal{L}}$, where $\bar{\mathcal{L}}$ is continuous from $H^{m}$ to $H^{1-m}$ and $H^{m}\times\mathcal{D}^{m}\ni(v,\phi)\mapsto[B(v\circ\phi^{-1})]\circ\phi\in L^{2}$ is differentiable. If the initial velocity $v_{0}$ has Sobolev regularity $m+k$ for $k\geq 1$, then so does the velocity $v_{t}$ for all $t$.
This result builds upon the knowledge acquired during recent years on $\mathcal{D}^{m}$, in particular the well-posedness of geodesics, the existence of shortest geodesics, and the rigorous derivation of the geodesic equation [MP10, BV17, GRR+23]. For smooth Riemannian metrics it can be derived from the right-invariance of the metric and the resulting Riemannian exponential map: Indeed, via the right-invariance, (higher order) differentiability of the exponential can be traded in for preservation of (higher order) Sobolev exponents [BHM25, Cor. 7.6], [BRU17, Sec. 5.2]. However, that argument does not exploit that the derivatives of the exponential (such as $D\exp_{\mathrm{id}}(v)$) need only be bounded in very specific directions (such as $Dv$) so that consequently the smoothness requirement on the metric can be weakened to the above. Our result confirms a conjecture of [MP10, Rem. 4.2] who could only show preservation of Sobolev regularity $m+k$ for $k>m+\frac{d}{2}$ in space dimension $d$ ([MP10, Thm. 4.1] is only stated for a particular class of smooth metrics, but the proof does not require the full smoothness).
@@
Finally note that LDDMM and the EPDiff equation are also frequently employed with alternative smoothing operators $\mathcal{R}$ such as convolution with a Gaussian. None of the shown analysis applies to these settings, since almost nothing is so far known about the geometry of the associated group of diffeomorphisms: Is it a manifold, and what (Banach) space $X$ is it modelled over? [^1] Note that some progress has been made for groups of diffeomorphisms that are flows of vector fields from non-Hilbert function spaces [NR19]. Is the right-invariant Riemannian metric induced by $\mathcal{L}=\mathcal{R}^{-1}$ smooth so that geodesics are well-defined? Our numerical analysis further requires an affirmative answer to the following additional questions: Is $X$ a Banach algebra so that one can make sense of the quadratic terms in (1)? Does an estimate of the form $\|(\mathcal{L}v)Dv\|_{(X^{+})^{\prime}}\leq\|v\|_{X}^{2}$ hold with $X^{+}=\{v\in X\,|\,Dv\in X^{d}\}$ and $(X^{+})^{\prime}$ its dual space? Do geodesics preserve the property $v_{t}\in X^{+}$ or even higher differentiability?
@@
Below we briefly introduce some notation employed throughout. We will work on the $d$-dimensional flat torus $\mathbb{T}^{d}$ (in applications $d\in\{2,3\}$), which will be identified with $0,1)^{d}$ with periodic boundary conditions. The single stroke norms $|\cdot|$ and $|\cdot|_{\infty}$ denote the Euclidean $\ell^{2}$- and the $\ell^{\infty}$-norm on finite-dimensional vectors or tensors. The identity matrix is $\mathrm{I}$, the identity operator on a function space is denoted by $\mathcal{I}$}, and the identity function by $\mathrm{id}$. If we write $\mathrm{id}:\mathbb{T}^{d}\to\mathbb{R}^{d}$, then we identify $\mathbb{T}^{d}$ with $0,1)^{d}$. The adjoint of a linear operator $P$ is written as $P^{*}$.
We will employ the following function spaces: $H^{l}(\mathbb{T}^{d})$ denotes the Sobolev space of scalar functions on the torus with square-integrable weak derivatives up to order $l$. For simplicity we will only work with integer orders; the special case $l=0$ refers to square-integrable Lebesgue functions. For $\mathbb{R}^{d}$-valued functions we use the notation $H^{l}(\mathbb{T}^{d};\mathbb{R}^{d})$. The notation $\|\cdot\|_{H^{l}}$ then refers to any one of the equivalent $H^{l}$-norms (for vector-valued functions simply taken componentwise). In contrast, $\vvvert\cdot\vvvert_{H^{m}}$ corresponds to one particular $H^{m}$-norm (the one chosen when defining a metric on the space of Sobolev diffeomorphisms). The corresponding dual spaces are denoted $H^{-l}(\mathbb{T}^{d})$ and $H^{-l}(\mathbb{T}^{d};\mathbb{R}^{d})$, respectively. Note that the diffeomorphisms will actually be functions in $H^{m}(\mathbb{T}^{d};\mathbb{T}^{d})$ (so domain and codomain are the torus); this simply means that the diffeomorphism is in $H^{m}$ if restricted to any simply connected neighbourhood and expressed in local coordinates of domain and co-domain (cf. [TAY02, § 4.3]; the local coordinates are obtained by the identification of $\mathbb{T}^{d}$ with $\mathbb{R}^{d}$ modulo the integer lattice). The Sobolev space of functions with $l$ essentially bounded derivatives is denoted by $W^{l,\infty}$ (domain and codomain are indicated as for $H^{l}$) with norm $\|\cdot\|_{W^{l,\infty}}$, where the special case $l=0$ indicates essentially bounded Lebesgue functions. $C^{n}$ denotes $n$ times continuously differentiable functions (again domain and codomain will be indicated as for $H^{l}$), with $C^{\infty}$ representing infinitely often differentiable functions. Finally, we will employ the Bochner space $L^{1}([0,1];H^{m}(\mathbb{T}^{d};\mathbb{R}^{d}))$ with norm $\|\cdot\|_{L^{1}([0,1];H^{m})}$ of absolutely integrable paths in $H^{m}(\mathbb{T}^{d};\mathbb{R}^{d})$.
@@ 2 The Riemannian manifold of Sobolev diffeomorphisms
It is well-known that $\mathcal{D}_{\mathrm{id}}^{m}$ is a differential ($C^{\infty}$) manifold modelled over the Hilbert space $H^{m}(\mathbb{T}^{d};\mathbb{R}^{d})$ [MP10]. In fact, this can be seen by interpreting $\mathcal{D}_{\mathrm{id}}^{m}$ as the quotient space $\widetilde{\mathcal{D}}_{\mathrm{id}}^{m}/\mathbb{Z}^{d}$ for
@@
Since $\mathcal{D}_{\mathrm{id}}^{m}$ (and thus also $\mathcal{D}^{m}$) is a group as well as a manifold, it behaves in some aspects like a Lie group and is often formally treated like one. However, inversion and left-multiplication are not smooth (not even differentiable), which is typically required for a Lie group (such groups have been termed half-Lie groups in [KMR15, MN18]). In fact, denoting the group of homeomorphisms and $C^{\infty}$-diffeomorphisms on the torus by $\mathrm{Hom}$ and $\mathrm{Diff}^{\infty}$, respectively, by [BOU72, III, Exerc. §4, ¶7] there cannot exist a Banach Lie group $G$ such that $\mathrm{Diff}^{\infty}\subset G\subset\mathrm{Hom}$ and the inclusions $\mathrm{Diff}^{\infty}\hookrightarrow G\hookrightarrow\mathrm{Hom}$ are continuous homomorphisms. In more detail, one can construct diffeomorphisms $\phi\in\mathrm{Diff}^{\infty}$ that are arbitrarily close to $\mathrm{id}$, but do not belong to any one-parameter subgroup of $\mathrm{Diff}^{\infty}$; thus they do not belong to a one-parameter subgroup of $G$ either. This, however, contradicts the local surjectivity of the Lie exponential map in Banach Lie groups.
@@
Let $(\!(\cdot,\cdot)\!)_{H^{m}}$ be an inner product of $H^{m}(\mathbb{T}^{d};\mathbb{R}^{d})$ with induced norm $\vvvert\cdot\vvvert_{H^{m}}$ equivalent to the standard $H^{m}$-norm, then
@@
is the induced right-invariant Riemannian Sobolev metric on $\mathcal{D}_{\mathrm{id}}^{m}$. [^2] Note that in contrast to most standard textbooks on Riemannian geometry we do not impose any smoothness requirements on a Riemannian metric, but simply take it to be a collection of inner products – one on each tangent space to the Riemannian manifold. Depending on the choice of $(\!(\cdot,\cdot)\!)_{H^{m}}$, (3) will of course entail some regularity of the map $\phi\mapsto g_{\phi}$, but this is ignored for now.
@@
The geodesic equation is the optimality condition for the minimization of $E$ and thus a second order ODE on the manifold $\mathcal{D}_{\mathrm{id}}^{m}$ (which is a PDE on $0,1]\times\mathbb{T}^{d}$). On honest Lie groups with right-invariant metric, the geodesic ODE is invariant under right action of the group on itself so that by Noether’s Theorem it can be reduced to a first order ODE. One can apply the same argument to formally derive such a first order geodesic equation on $\mathcal{D}_{\mathrm{id}}^{m}$: Let $\phi$ be a geodesic path with velocity $v$ and let $\eta\in C^{\infty}([0,1]\times\mathbb{T}^{d})$ be an infinitesimal perturbation with $\eta_{0}=\eta_{1}=0$. Defining $\phi^{\varepsilon}_{t}=(\mathrm{id}+\varepsilon\eta_{t})\circ\phi_{t}$, its velocity is given by $v^{\varepsilon}_{t}=\dot{\phi}^{\varepsilon}_{t}\circ(\phi^{\varepsilon}_{t})^{-1}=\left[(\mathrm{I}+\varepsilon D\eta_{t})v_{t}+\varepsilon\dot{\eta}_{t}\right]\circ(\mathrm{id}+\varepsilon\eta_{t})^{-1}$. Assuming sufficient differentiability, the optimality conditions read
@@
This first order ODE is known as the EPDiff or Euler–Poincaré equation [TY15]; its integral yields $\rho$ or equivalently $v$, from which then the geodesic path $\phi$ can be calculated as the flow. Note that the derivation was merely formal since along the way we illegally differentiated the group product of $\mathcal{D}_{\mathrm{id}}^{m}$ (e.g. $\mathrm{Ad}_{\phi}$ with $\phi\in\mathcal{D}_{\mathrm{id}}^{m}$ maps $H^{m}(\mathbb{T}^{d};\mathbb{R}^{d})$ only into $H^{m-1}(\mathbb{T}^{d};\mathbb{R}^{d})$). As a consequence, due to $v_{t}\in H^{m}(\mathbb{T}^{d};\mathbb{R}^{d})$ and $\rho_{t}\in H^{-m}(\mathbb{T}^{d};\mathbb{R}^{d})$, the right-hand side of (6) only lies in $H^{-m-1}(\mathbb{T}^{d};\mathbb{R}^{d})$, and the equation may not be well-defined.
@@
    paths of finite energy in $\mathcal{D}_{\mathrm{id}}^{m}$ exist between any two elements of $\mathcal{D}_{\mathrm{id}}^{m}$ (since $\mathcal{D}^{m}$ is open and locally connected in $H^{m}(\mathbb{T}^{d};\mathbb{R}^{d})$, its connected components, in particular $\mathcal{D}_{\mathrm{id}}^{m}$, are open; now connected open subsets of normed spaces are polygonally connected, and a polygon in $\mathcal{D}_{\mathrm{id}}^{m}$, consisting of linear interpolation segments $t\mapsto t\phi_{1}+(1-t)\phi_{0}$, has finite path energy),
@@
    for the inner product $(\!(v,w)\!)_{H^{m}}=\langle\mathcal{L}v,w\rangle$ with $\mathcal{L}=(1-\Delta)^{m}$ or a more general differential operator of order $2m$ with smooth coefficients the metric is smooth [MP10, Thm. 4.1] (this even holds for any elliptic invertible operator $\mathcal{L}\in\mathrm{OPS}_{1,0}^{2m}$ of order $2m$ [BBC+20, Thm. 4.16 & Thm. 5.1], see [EK14, BEK15] for the Fourier multiplier case), thus the geodesic ODE is locally uniquely solvable by classical differential geometry and ODE theory [KLI95, § 1.6], its solution depends smoothly on the initial conditions, and by the Inverse Function Theorem the Riemannian exponential map is a local diffeomorphism,